Ethane is burned with 50% excess air. the percentage conversionof the ethane is 90%. from the ethane that is converted , 25% react to form CO and the rest react to form CARBONDIOXIDE.
what is the molar composition of the stack gas on a dry basis? what is the mole ratio of water to dry stack gas
1
Expert's answer
2020-03-23T10:13:37-0400

n0 - initial

n - final

2C2​H6​+7O2​→4CO2​+6H2​O

2C2​H6​+5O2​→4CO+6H2​O

Let n0(C2​H6​) be 1 mol then

n0(O2​) = 1/2*7*1.5 = 5.25 mol

n(C2​H6​) = 0.1 mol

n(O2) = 5.25 - 1*0.9*0.25/2*5 - 1*0.9*0.75/2*7 = 2.325 mol

n(CO2) = 1*0.9*0.75/2*4 = 1.35 mol

n(CO) = 1*0.9*0.25/2*4 = 0.45 mol

n(H2O) = 1*0.9*0.75/2*6 + 1*0.9*0.25/2*6 = 2.7 mol

the molar composition of the stack gas on a dry basis:

C2​H6 0.1/(0.1+2.325+1.35+0.45)*100 = 2.4%

O2 2.325/(0.1+2.325+1.35+0.45)*100 = 55.0%

CO2 1.35/(0.1+2.325+1.35+0.45)*100 = 32.0%

CO 0.45/(0.1+2.325+1.35+0.45)*100 = 10.6%

the mole ratio of water to dry stack gas:

2.7/(0.1+2.325+1.35+0.45) = 0.639